OpenStudy (anonymous):

I have the answer can someone just check to see if I'm right. WHich of the following is f '(x) for f(x)=(ln4)^x

zepdrix (zepdrix):

So you recall that \(\large e^x\) gives us \(\large e^x\) when we differentiate it, yes? Whenever we differentiate an exponential function that is NOT BASE e, we have to multiply by the natural log of the base. Example:\[\large y=2^x \qquad \rightarrow \qquad y'=2^x(\ln 2)\] We have in this problem,\[\large y=(\ln 4)^x \qquad \rightarrow \qquad y'=(\ln 4)^x(\ln(\ln4))\]It's not pretty, I know :) lol

OpenStudy (anonymous):

ok gotit that is the answer that I had to

zepdrix (zepdrix):

cool c: